# Aluminum nuclear demagnetization refrigerator for powerful continuous cooling

### v1

@@ Abstract
Many laboratories routinely cool samples to 10 mK, but relatively few can cool condensed matter below 1 mK. Easy access to the microkelvin range would propel fields such as quantum sensors and quantum materials. Such temperatures are achieved with adiabatic nuclear demagnetization. Existing nuclear demagnetization refrigerators (NDR) are “single-shot”, and the recycling time is incompatible with proposed sub-mK experiments. Furthermore, a high cooling power is required to overcome the excess heat load of order nW on NDR pre-cooled by cryogen-free dilution refrigerators. We report the performance of an aluminum NDR designed for powerful cooling when part of a dual stage continuous NDR (CNDR). Its thermal resistance is minimized to maximize the cycling rate of the CNDR and consequently its cooling power. At the same time, its susceptibility to eddy current heating is minimized. A CNDR based on two of the aluminum NDR presented here would have a cooling power of approximately 40 nW at 560 $\mu$K.
@@ I Introduction
Cooling condensed matter well below the $\approx 5$ mK limit of dilution refrigeration was motivated in part by the rich physics generated by the internal degrees of freedom of superfluid 3He. Studies of this system continue to generate fascinating new results [1]. Today, a broad spectrum of research is conducted near 1 mK and far below. These include quantum criticality in strongly correlated electron systems [2, 3], nanomechanical resonators [4], amorphous solids [5, 6, 7], electronic transport in nanostructures [8], 2D electron gases [9] and dark matter searches [10, 11].
@@
Continuous cooling below 1 mK using two NDR has been proposed [28, 29]. The NDR stages may be arranged in a serial or parallel configuration [30]. One design for continuous NDR (CNDR) relies on PrNi5 refrigerant, with a projected cooling power of 10 nW at 1 mK [31]. Advantages of this refrigerant include the enhancement of the magnetic field experienced by the nuclei to $\approx 10$ times the applied one, the small Korringa constant $<\kappa\approx 1$ mK sec yielding rapid thermal equilibration of nuclei and electrons, and a high electrical resistance yielding low eddy current heating [32, 33]. These features of PrNi5 are balanced by its fragility [31], difficulties in its synthesis that may lead to magnetic losses [32], its high cost and limited availability, the minimum temperature $\approx 0.4$ mK that is limited by magnetic ordering of the nuclei [34], and the low thermal conductance of the alloy. The thermal resistance of a PrNi5 stage can be improved by soldering Cu or Ag conductors to the sides of the PrNi5 rods. However, the 29 mT critical field of indium solder [20] and the similarly strong internal magnetic field of PrNi5 [34] limit the final applied field.
@@
We report the performance of a new aluminum nuclear demagnetization refrigerator. The nuclear stage is composed of a bundle of 4N pure aluminum wires, which simultaneously optimizes the thermal conductance along the nuclear stage and the eddy current heating. The system operates at a relatively low maximum magnetic field of $\approx 3$ T. This allows efficient reduction of stray fields so that samples and thermometers sensitive to magnetic fields can be used. In particular, a commercially available SQUID-based noise thermometer was used in the present work. Furthermore, the size of the two main coils of the future CNDR will be relatively small and the interaction between the coils will be minimal.
@@ II Design
Figure 1: Schematic of the aluminum wire demagnetization stage and magnets mounted on the dilution pre-cooler. The Cu plate is connected to the MXC plate via a heat switch (not shown).
The nuclear stage consisted of a bundle of 280 aluminum wires of 1 mm diameter and 4 N purity. It was centered inside the main coil with nylon screws. The wires were welded at the top and at the bottom of the bundle. The weld at the top provided a platform for a ruthenium oxide resistance thermometer and a heater H1. The weld at the bottom was gold plated and pressed to a gold plated copper block, using the technique described in [37]. A Magnicon MFFT-1 magnetic field fluctuation thermometer was attached to the copper block. The copper block was also connected to the heat switch described in [35] using a series of copper pieces, including a copper plate on which we mounted a carbon resistance thermometer R54 and a heater H3. The copper plate was attached to the mixing chamber plate with carbon rods. The other terminal of the heat switch was bolted to the top of the mixing chamber plate, whose temperature was monitored with another MFFT-1 noise thermometer.
@@ III Performance
We measured temperatures below 800 $\mu$K using the MFFT attached to the Cu block. In one run, we precooled to $T_{\mathrm{MFFT}}=7.7$ mK while applying $B_{\mathrm{main}}=2.84$ T using the main coil. We then demagnetized the nuclear stage to $B_{\mathrm{main}}=140$ mT (Fig. 2). Ten minutes after reaching the final magnetic field, the temperature stabilized at $T_{\mathrm{MFFT}}=880$ $\mu$K. We then turned off the pulse tube and observed further cooling. Ten minutes later, just before it became necessary to turn the pulse tube back on, $T_{\mathrm{MFFT}}$ had reached 790 $\mu$K and was continuing to decrease. In the absence of losses, we would expect a nuclear temperature $T_{\mathrm{ideal}}=385$ $\mu$K due to the 20-fold reduction in $B_{\mathrm{main}}$. We believe that the sensitivity of $T_{\mathrm{MFFT}}$ to pulse tube vibrations could be decreased by decreasing the contact resistance between the MFFT copper body and the copper block to which it is screwed. The MFFT body is not gold plated and its tightening torque is limited by the strength of the highly pure copper from which it is constructed. The small temperature gradient along the nuclear stage and the temperature difference between the electrons and nuclei of the nuclear stage are estimated below. Furthermore, losses in the demagnetization process elevate the minimum nuclear temperature above $T_{\mathrm{ideal}}$. The losses due to vibrations and eddy current heating are quantified below.
Figure 2: The temperature of the nuclear stage was measured with a noise thermometer [39] screwed to a Au-plated Cu extension of the Cu block (see text), as shown in the inset.
@@
The heat capacity measurements shown in the main panel of Fig. 3 were obtained from pulses like those in the inset by dividing the deposited energy by the increase in equilibrium temperature. The size of the temperature step was typically 10% of the average temperature. The curves correspond to the theoretical nuclear heat capacity [40] of 2.0 mol of aluminum in the magnetic fields specified in the legend, which correspond to the applied field at the center of the main coil. The theoretical curves are in good agreement with measurements at the highest fields, so that 2.0 mol can be considered the effective quantity of aluminum in the nuclear stage. This value is close to the 2.2 mol of aluminum inside the 10 cm long main coil, and the 10% difference may be explained by the inhomogeneity of the magnetic field. At low values of $B/T$ we observed an excess heat capacity that cannot be explained by the electronic contribution in these 2 mol of aluminum. The copper in the thermal link between the heat switch and the nuclear stage also contributes to the measured heat capacity and may be responsible for the observed excess. An anomalous heat capacity in Cu was also observed in [41] and attributed to electric field gradients in the Cu due to lattice defects.
Figure 3: Heat capacity of the nuclear stage (points) compared with the theoretical nuclear heat capacity at the specified magnetic fields (curves). Inset: Response of the nuclear stage to a heat pulse (see main text).
@@
In order to determine the residual heat load in the low field limit, we opened the heat switch and used it as a known thermal resistance. In particular, we first applied heater powers to the nuclear stage that were much greater than the estimated parasitic heat load. This allowed us to verify that the open heat switch had the same temperature dependent thermal resistance as reported in [35]. We then turned off the applied heat and allowed the system to stabilize. The heat load corresponding to the equilibrium temperature of the nuclear stage was obtained by consulting the measured heat flow as a function of the temperature of the hot side of the switch reported in [35]. The temperature of the cold side of the switch was close to the base temperature of the dilution refrigerator in both cases and was therefore negligible. The upper inset of Fig. 5 shows the time dependence of the heat load at $B_{\mathrm{main}}=28$ mT. A portion of the time dependent heat load may come from the aluminum, which is known to release comparable amounts of heat per unit mass even in samples that are much more pure than the aluminum used here [44].
@@
The contribution to eddy current heating that would be present in the absence of vibrations was quantified by applying an oscillating magnetic field to the nuclear stage. In particular, the stage temperature was regulated at 130 mK with the heat switch open at a particular static field $B_{\mathrm{main}}$. We then oscillated $B_{\mathrm{main}}$ with amplitude 28 mT around the initial static field. The resulting decrease in the heater power required to maintain the stage at 130 mK corresponded to the eddy current heating. These results are plotted in the inset of Fig. 6. As expected, we observed a quadratic dependence of eddy current heating on $\partial B/\partial t$. The corresponding heating coefficient is plotted as a function of average $B_{\mathrm{main}}$ in the main panel of Fig. 6. We believe that this coefficient decreases with average $B_{\mathrm{main}}$ due to magnetoresistance of the aluminum [45]. While magnetoresistance of copper has a large influence on eddy current heating in systems such as the one described in [20], where electrical resistance increases by a factor of over 20 in 6 tesla, the magnetoresistance of aluminum is different [46]. The maximum demagnetization rate in the present work was 5.1 T/h, so that this contribution to the heat load was always below $\approx 20$ nW.
@@ IV Conclusion
We therefore expect a cooling power of at least 40 nW at 560 $\mu$K. The $\approx 10$ nW low field parasitic load measured in the present work would yield a negligible 40 $\mu$K decoupling of the electrons from the nuclei [48]. We therefore anticipate that a CNDR based on Al stages of the type described here will maintain temperatures well below 1 mK indefinitely.

### v2

@@ Abstract
Many laboratories routinely cool samples to 10 mK, but relatively few can cool condensed matter below 1 mK. Easy access to the microkelvin range would propel fields such as quantum sensors and quantum materials. Such temperatures are achieved with adiabatic nuclear demagnetization. Existing nuclear demagnetization refrigerators (NDR) are “single-shot”, and the recycling time is incompatible with some sub-mK experiments. Furthermore, a high cooling power is required to overcome the excess heat load of order nW on NDR pre-cooled by cryogen-free dilution refrigerators. We report the performance of an aluminum NDR designed for powerful cooling when part of a dual stage continuous NDR (CNDR). Its thermal resistance is minimized to maximize the cycling rate of the CNDR and consequently its cooling power. At the same time, its susceptibility to eddy current heating is minimized. A CNDR based on two of the aluminum NDR presented here would achieve a cooling power of approximately 40 nW at 560 $\mu$K less than six days after cooling from room temperature, with a small offset in electronic temperature that decreases as the time-dependent heat load decays.
@@ I Introduction
Cooling condensed matter well below the $\approx 5$ mK limit of dilution refrigeration was motivated in part by the rich physics generated by the internal degrees of freedom of superfluid 3He. Studies of this system continue to generate fascinating new results [1, 2, 3, 4, 5, 6]. Today, a broad spectrum of research is conducted near 1 mK and far below. These include quantum criticality in strongly correlated electron systems [7, 8], nanomechanical resonators [9], amorphous solids [10, 11, 12], electronic transport in nanostructures [13], 2D electron gases [14] and dark matter searches [15, 16].
@@
Continuous cooling below 1 mK using two NDR has been proposed [33, 34]. The NDR stages may be arranged in a serial or parallel configuration [35]. One design for continuous NDR (CNDR) relies on PrNi5 refrigerant, with a projected cooling power of 10 nW at 1 mK [36]. Advantages of this refrigerant include the enhancement of the magnetic field experienced by the nuclei to $\approx 10$ times the applied one, the small Korringa constant $\kappa\lesssim 1$ mK sec (compared with 1.80, 1.27 and 1.09 K sec for 27Al, 63Cu and 65Cu, respectively) yielding rapid thermal equilibration of nuclei and electrons, and a high electrical resistance yielding low eddy current heating [37, 38]. These features of PrNi5 are balanced by its fragility [36], difficulties in its synthesis that may lead to magnetic losses [37], its high cost and limited availability, the minimum temperature $\approx 0.4$ mK that is limited by magnetic ordering of the nuclei [39], and the low thermal conductance of the alloy. The thermal resistance of a PrNi5 stage can be improved by soldering Cu or Ag conductors to the sides of the PrNi5 rods. However, the 29 mT critical field of indium solder [25] and the similarly strong internal magnetic field of PrNi5 [39] limit the final applied field.
@@
We report the performance of a new aluminum nuclear demagnetization refrigerator. The nuclear stage is composed of a bundle of 4N pure aluminum wires, which simultaneously optimizes the thermal conductance along the nuclear stage and the eddy current heating. The system operates at a relatively low maximum magnetic field of $\approx 3$ T. This allows efficient reduction of stray fields so that samples and thermometers sensitive to magnetic fields can be used. In particular, a SQUID-based magnetic field fluctuation thermometer (MFFT) [42, 43] was used in the present work. Furthermore, the size of the two main coils of the future CNDR will be relatively small and the interaction between the coils will be minimal.
@@ II Design
Figure 1: Schematic of the aluminum wire demagnetization stage and magnets mounted on the dilution pre-cooler. The Cu plate is connected to the MXC plate via a heat switch (not shown). Metal film heaters H1 and H3 are placed at the top and bottom of the nuclear stage, a ruthenium oxide thermometer is at the top of the stage, and a carbon resistor R54 and magnetic field fluctuation thermometer are at the bottom of the stage. Nylon screws are used to center the Al wire bundle.
The nuclear stage consisted of a bundle of 280 aluminum wires of 1 mm diameter and 4 N purity. It was centered inside the main coil with nylon screws. The wires were welded at the top and at the bottom of the bundle. The weld at the top provided a platform for a ruthenium oxide resistance thermometer and a heater H1. The weld at the bottom was gold plated and pressed to a gold plated copper block, using the technique described in [44]. The MFFT was attached to the copper block. The copper block was also connected to the heat switch described in [40] using a series of copper pieces, including a copper plate on which we mounted a carbon resistance thermometer R54 and a heater H3. The copper plate was attached to the mixing chamber plate with carbon rods. The other terminal of the heat switch was bolted to the top of the mixing chamber plate, whose temperature was monitored with another MFFT.
@@ III Performance
We measured temperatures below 1 mK using the MFFT attached to the Cu block less than six days after starting to cool from room temperature. In one run, we precooled to $T_{\mathrm{MFFT}}=7.7$ mK while applying $B_{\mathrm{main}}=2.84$ T using the main coil. We then demagnetized the nuclear stage to $B_{\mathrm{main}}=140$ mT (Fig. 2). Ten minutes after reaching the final magnetic field, the temperature stabilized at $T_{\mathrm{MFFT}}=880$ $\mu$K. We then turned off the pulse tube and observed further cooling. Ten minutes later, just before it became necessary to turn the pulse tube back on, $T_{\mathrm{MFFT}}$ had reached 790 $\mu$K and was continuing to decrease. This test was carried out at the end of the “cooldown”, four weeks after starting to cool from room temperature. While we expect some improvement in performance with time due to the decay of the time-dependent heat leak (see below), the delay in carrying out this test involving the pulse-tube shutdown was primarily due to other practical issues. In the absence of losses, we would expect a nuclear temperature $T_{\mathrm{ideal}}=385$ $\mu$K due to the 20-fold reduction in $B_{\mathrm{main}}$. We believe that the sensitivity of $T_{\mathrm{MFFT}}$ to pulse tube vibrations could be decreased by decreasing the contact resistance between the MFFT copper body and the copper block to which it is screwed. The MFFT body is not gold plated and its tightening torque is limited by the strength of the highly pure copper from which it is constructed. The small temperature gradient along the nuclear stage and the temperature difference between the electrons and nuclei of the nuclear stage are estimated below. Furthermore, losses in the demagnetization process elevate the minimum nuclear temperature above $T_{\mathrm{ideal}}$. The losses due to vibrations and eddy current heating are quantified below.
Figure 2: Electronic nuclear stage temperatures below 1 mK were measured with the noise thermometer (MFFT) less than six days after starting to cool from room temperature. The temperature record at the end of a demagnetization made three weeks later (main panel) showed that pulse tube operation was causing an offset in the MFFT temperature. The inset shows how the MFFT was screwed to a Au-plated Cu extension of the Cu block (see also Fig. 1).
@@
The heat capacity measurements shown in the main panel of Fig. 3 were obtained from pulses like those in the inset by dividing the deposited energy by the increase in equilibrium temperature. The size of the temperature step was typically 10% of the average temperature. The curves correspond to the theoretical nuclear heat capacity [46] of 2.0 mol of aluminum in the magnetic fields specified in the legend, which correspond to the applied field at the center of the main coil. The theoretical curves are in good agreement with measurements at the highest fields, so that 2.0 mol can be considered the effective quantity of aluminum in the nuclear stage. This value is close to the 2.2 mol of aluminum inside the 10 cm long main coil, and the 10% difference may be explained by the inhomogeneity of the magnetic field. At low values of $B/T$ we observed an excess heat capacity that cannot be explained by the electronic contribution in these 2 mol of aluminum. The copper in the thermal link between the heat switch and the nuclear stage also contributes to the measured heat capacity and may be responsible for the observed excess. An anomalous heat capacity in Cu was also observed in [47] and attributed to electric field gradients in the Cu due to lattice defects. While the excess heat capacity may temporarily load the proposed CNDR, we nonetheless expect the CNDR performance discussed in the conclusion since we achieved $T_{\mathrm{MFFT}}<1$ mK less than six days after cooling from room temperature.
Figure 3: The heat capacity of the nuclear stage (points) at different main coil currents was determined from the increase in equilibrium temperature following applied heat pulses. The measured heat capacity at high magnetic fields agreed with the theoretical nuclear heat capacity (curves) of 2 moles of aluminum at the specified fields, which were the maximum magnetic fields in the main coil corresponding to the currents. Inset: Response of the nuclear stage to a heat pulse demonstrating the short equilibration time (see main text).
@@
In order to determine the residual heat load in the low field limit, we opened the heat switch and used it as a known thermal resistance. In particular, we first applied heater powers to the nuclear stage that were much greater than the estimated parasitic heat load. This allowed us to verify that the open heat switch had the same temperature dependent thermal resistance as reported in [40]. We then turned off the applied heat and allowed the system to stabilize. The heat load corresponding to the equilibrium temperature of the nuclear stage was obtained by consulting the measured heat flow as a function of the temperature of the hot side of the switch reported in [40]. The temperature of the cold side of the switch was close to the base temperature of the dilution refrigerator in both cases and was therefore negligible. The upper inset of Fig. 5 shows the time dependence of the heat load at $B_{\mathrm{main}}=28$ mT. Time-dependent heat leaks of a similar initial magnitude decaying approximately exponentially with a time constant of one week were previously observed in Cu and PrNi5 based NDR [50, 51]. A portion of our time dependent heat load may come from the aluminum, which is known to release comparable amounts of heat per unit mass even in samples that are much more pure than the aluminum used here. This heat release was explained in terms of relaxation of tunneling systems [52]. Furthermore, molecular hydrogen may become trapped in Cu and release heat when undergoing an ortho-para conversation at cryogenic temperatures. The time-dependent heat leak of our NDR may therefore be improved by annealing Cu parts that were not heat treated [38].
@@
The contribution to eddy current heating that would be present in the absence of vibrations was quantified by applying an oscillating magnetic field to the nuclear stage. In particular, the stage temperature was regulated at 130 mK with the heat switch open at a particular static field $B_{\mathrm{main}}$. We then oscillated $B_{\mathrm{main}}$ with amplitude 28 mT around the initial static field. The resulting decrease in the heater power required to maintain the stage at 130 mK corresponded to the eddy current heating. These results are plotted in the inset of Fig. 6. As expected, we observed a quadratic dependence of eddy current heating on $\partial B/\partial t$. The corresponding heating coefficient is plotted as a function of average $B_{\mathrm{main}}$ in the main panel of Fig. 6. We believe that this coefficient decreases with average $B_{\mathrm{main}}$ due to magnetoresistance of the aluminum, which, to our knowledge, has not been measured in the low mK range [53]. While magnetoresistance of copper has a large influence on eddy current heating in systems such as the one described in [25], where electrical resistance increases by a factor of over 20 in 6 tesla, the magnetoresistance of aluminum is different [54]. The maximum demagnetization rate in the present work was 5.1 T/h, so that this contribution to the heat load was always below $\approx 20$ nW. In another cooldown, we characterized the non-adiabaticity of the NDR by determining the equilibrium value of $B_{\mathrm{main}}/T_{\mathrm{MFFT}}$ in a temperature range where decoupling of the thermometer is negligible. In particular, we started at $T_{\mathrm{MFFT}}=6.73$ mK, yielding $B_{\mathrm{main}}/T_{\mathrm{MFFT}}=0.42$ T/mK at the initial $B_{\mathrm{main}}=2.84$ T. We demagnetized at 1.6 T/h, pausing for 1 hour after each 0.28 T decrease in $B_{\mathrm{main}}$. When $T_{\mathrm{MFFT}}$ reached 3.0 mK, $B_{\mathrm{main}}/T_{\mathrm{MFFT}}$ had decreased to 0.38 T/mK.
@@ IV Conclusion
Quickly achieving $T<1$ mK, starting from room temperature, may be important in some applications. In a subsequent cooldown of the same NDR, after adjusting the centering screws (Fig. 1), we achieved somewhat higher absorption capacity less than six days after starting to cool from room temperature, measuring $T_{\mathrm{MFFT}}=1.85$ mK at $B_{\mathrm{main}}=568$ mT.
@@
We therefore expect a cooling power of at least 40 nW at 560 $\mu$K less than six days after cooling from room temperature. The low field heat leak measured in the present work, which decreased from 30 to 7 nW, would yield a small decoupling of the electrons from the nuclei ranging from 115 to 27 $\mu$K [56]. We therefore anticipate that a CNDR based on Al stages of the type described here will maintain temperatures well below 1 mK indefinitely.